I am trying to help my daughter with homework. Since it has been many years I have forgotten how to do. How to i anwer this ...Seperate $89 into two parts so that the second part is $4less than twice the first part.
Answer by stanbon(75887)

Separate $89 into two parts so that the second part is $4 less than twice the first part.
-------------------
Let the 1st part be "x"
Then 2nd part is "2x-4"
EQUATION:
1st + 2nd = 89
x+2x-4=89
3x-4=89
3x=93
x=31 (This is the "1st part")
2x-4=62-58 (This is the "2nd part")
Cheers,
Stan H.